OpenStudy (anonymous):

hey guys, sal here i was wondering if you fellas could help me with these functions?

OpenStudy (anonymous):

http://vvcap.net/db/9KWxL68W2p-0wUCsfgeS.htp

OpenStudy (anonymous):

F(3)= 3x+1?? for the first one?

OpenStudy (anonymous):

ok i will show work for the first one .follow the same for the second one since twwo points are given (3,1) and (7,13) find the slope \[m=\frac{\Delta y}{\Delta x}\] m=(13-1)/(7-3)=12/4=3 using point slope form y-y0=m(x-x0) y-1=3(x-3) y-1=3x-9 y=3x-8 f(x)=y=3x-8

OpenStudy (anonymous):

Given that f is a linear function : f(x) = mx + b, and f(h) = k, f(m) = n for instance, then we have \[\large y - k = \frac{n-k}{m-h}(x - h)\] \[\large y = \frac{n-k}{m-h}x + h\left( \frac{k-n}{m-h} \right)+k\]

OpenStudy (anonymous):

awesome guys, thanks!